OpenStudy (anonymous):

!!!! High importance! Will fan/medal, just need to check answers! Nitrogen can ionically bond with an unknown element X from group 2. How many ions of element X are required to create a stable ionic compound with nitrogen? A. 2 B. 3 C. 4 D. There's not enough information. My answer is B, am I correct?

OpenStudy (cuanchi):

All alkaline earth metals are competent to producing their appropriate nitrides in reaction with nitrogen with a required vast amount of heat. The basic structure of reactions with an alkaline earth metal and nitrogen are as follows: 3M(s)+N2(g)−→M3N2(s) 3M(s)+N2(g)→M3N2(s) (M = Be, Mg, Ca, Sr, or Ba) Nitrides of the alkaline earth metals are ionic and solid in nature except for beryllium nitride (Be3N2Be3N2), which is covalent and unpredictable. Decompositions of the alkaline earth metal nitrides occur upon heating: M3N2(s)−→3M(s)+N2(g) M3N2(s)→3M(s)+N2(g) (M = Be, Mg, Ca, Sr, or Ba) Nitrides of alkaline earth metals may also react with water producing ammonia: M3N2 (s) + 6H2O (l) → 3M(OH)2 (aq) + 2NH3 (g) (M = Be, Mg, Ca, Sr, or Ba) http://chemwiki.ucdavis.edu/Core/Inorganic_Chemistry/Descriptive_Chemistry/Main_Group_Reactions/Reactions_of_Main_Group_Elements_with_Nitrogen